how do I solve for y: 3x+y=3
Answer by kekegeter (97)

3x+y=3
(3x+y)+(-3x)=3+(-3x)
(3x+y)-3x=3-3x
3x+y-3x-3x+3
3x-3x+y=-3x+3
y=-3x+3
All done!